### v1

# Kinetic coupling and dark matter on the lattice K. Farakosa, V.Iconomoua and G. Koutsoumbasa

Abstract
We study a system coupled to scalar fields. Initially the model is studied in a novel continuum formulation and study of the appropriate diagonalizations is performed. Three models are examined, in two of which the scalar field couples with both gauge fields while in the third one the scalar field couples only to the dark photon. The model is then treated on a space-time lattice. We determine the phase diagram for various values of the kinetic coupling parameter. Then there follows the determination of masses for the scalar fields and the massive gauge fields, as well as the fine structure constants for the massless gauge fields.
1 Introduction
The biggest part of matter in the Universe is of unknown nature. Our current knowledge suggests that no more than of the mass-energy of the Universe can be attributed to ordinary matter, while nearly corresponds to this unknown form of matter known as Dark Matter [1]. The remaining correspond to Dark Energy, which is a hypothetical substance driving the accelerating expansion of the Universe. The nature of dark matter and dark energy are among the greatest puzzles of modern cosmology and astrophysics. Dark matter has only been observed through its gravitational effects with ordinary matter and its fundamental properties are still a mystery. The existence of dark matter suggests the existence of new physics.
Now we present indications about dark matter coming mainly from gravitational observations. Among the first indications for the existence of dark matter came from the consideration of the motion of galaxies within the Coma Cluster [2]. The mass of this cluster has been estimated in two ways, one based on total luminosity and another based on the virial theorem. The inconsistency of the two results has given rise to the hypothesis that matter had also some dark component which could not be detected by the telescopes [3, 4].
A rotation curve of stars in galaxies is a measurement of the orbital velocity of stars around the center of the galaxy as a function of their distance from the galactic center. The rotational velocity was expected to decrease, however they were measured to be essentially constant at large distances [5], so that there should be additional, not visible, i.e. dark, matter, to account for this behaviour.
Gravitational lensing offers extra support to the Dark Matter hypothesis. Study of the images of distant galaxies may provide clues about the distribution of Dark Matter in space [6, 7]. Intra-cluster gas within galaxy clusters may produce X rays, whose lensing results may yield useful information about the distribution of dark matter [8, 9].
The study of the Bullet Cluster [10, 11] provides some of the most compelling observational evidence for the existence of dark matter to date. Further evidence may be found in [12], describing MACS J0025.4-1222 cluster collisions. Evidence in favour of dark matter is also provided by the fluctuations of the Cosmic Microwave Background, for example, the Sachs-Wolfe effect [13]. Finally, the Universe appears very inhomogeneous today, however CMB was very smooth in the early Universe. It is difficult to explain the structure formation without the presence of dark matter, since the magnitude of fluctuations would be too small. Interesting approaches to dark matter in stars have been proposed in [14]. For a very partial, but characteristic, sample of references from the experimental/phenomenological side, on millicharged dark-matter searches, one may consult [15] and references therein.
Although unrelated to dark matter problem, the idea of a new force under which Standard Model particles are neutral emerged in 1986 by Holdom [16]. Several aspects of this idea have since appeared in the literature in the context of grand unified theories [17], or dark matter models [18] or string phenomenology [19]. There have been proposals for experimental detection of additional s [20], as well as experimental trials to detect dark photons [21]. One may find works on the kinetic coupling and its relation to astrophysics [22]. This hidden force has provided new theoretical and experimental directions in the search for particle dark matter. The extension of the Standard Model by the addition of the dark mediated by a dark photon, also known as “hidden photon”, was originally proposed in [9], [23]. The models can contain a massless dark photon, with an unbroken In a theory that includes two groups, a kinetic mixing term may arise between the two gauge bosons [16]. The dark sector is considered as a sector interacting gravitationally, as well as weakly, with the standard model sector, for example through kinetic mixing of the dark and the visible gauge field. The gravitational interaction alone is not enough to produce effects which could serve to detect dark sector particles in the laboratory.
The kinetic mixing between two gauge groups establishes a communication channel between the dark and the visible gauge sector. The dark photon acting as a mediator between the dark and visible sector is known as the “vector portal”, since it is a spin-1 particle. However, other viable portals have been proposed mediated by dark particles of different spins, such as the “scalar portal” mediated by scalar DM particle or “neutrino portal” mediated by a dark fermionic field like a right-handed neutrino [24].
The dark photon may be chosen to couple to either the electromagnetic or the hyper-charge of the standard model [25], [26]. The choice of the appropriate mixing is related to the energy of the processes under examination: above the electroweak symmetry breaking scale, it is reasonable to couple to the hypercharge, while below that scale, the coupling to the physical photon is relevant.
Among the special features of the models is that the corresponding field strength is not just covariant, as is the case with the non-abelian models, but it is also gauge invariant. This feature, along with the presence of various factors, opens up the way to include mixed terms in the Lagrangian, consisting of products of different s. These considerations arose within the treatment of models of dark matter. If the s are embedded in a grand unification group, the mixing will be zero at the unification scale, but kinetic coupling may arise through renormalisation group effects at lower scales, where the original group has broken down.
There have been several approaches to the study of models with additional gauge groups. Kinetic mixing is a recurring theme in these studies. An early one may be found in [26]. A nice description of the relevant models may be found in [25]. This ends up to a new set of decoupled gauge fields. A central issue is the procedure to decouple the two gauge fields, which are coupled through the kinetic coupling, as well as through the symmetry breaking process. It turns out that the transformation is not orthogonal. A by now standard practice is to leave the dark gauge field intact and transform from the visible gauge field to a new one, which contains a small admixture of the dark field. In view of the lattice formulation we propose a different method, which is not orthogonal in the continuum either, however it remains orthogonal in the lattice formalism, the departure from orthogonality been reflected in just a redefinition of lattice parameters. This facilitates the orthogonalization procedure, as will be shown. We describe the two methods of decoupling in the continuum, as well as the decoupling in the lattice formulation of the model. We note that the massive and massless gauge boson cases have been studied separately at tree level, the mass generation mechanism being left unspecified. In this work we employ a higgs field to generate masses for one of the two gauge fields, the other one remaining massless. A unified description of both the massless and the massive gauge bosons follows from this formulation.
This model has been studied by tree level calculations and it is an important issue to proceed to the full quantum model via non-perturbative techniques, such as lattice Monte Carlo simulations, since symmetry breaking is an essentially non-perturbative effect. To the best of our knowledge, the numerical investigations of such models have not been done in the past, therefore we developed efficient algorithmic software which is applicable on massively parallel computing systems for the Monte Carlo generation of appropriate configurations and the numerical calculation of the related field-theoretical correlation functions. The point is that, to identify the phases of the resulting model, one should calculate numerically highly non-local quantities, such as the Wilson loops, which will give the potentials felt by heavy quarks in each phase.
This work starts with a novel formulation of the model in the continuum, section 2. In particular we present the gauge fields in the presence of the kinetic mixing and present the diagonalization procedure, when matter fields are either missing or present; we identify the massless and massive fields (with their mass) in terms of the original visible and dark fields at the tree level. The results are specified for each of three models (a)model A, where the scalar field charges equal (b) model B, where the charges equal and (c) model C, with In models A and C both gauge fields are coupled with the scalar field, while in model B the scalar field does not feel the visible gauge field, but only the dark gauge field. Next we present the by now standard formulation of the problem. This procedure is described e.g. in [25]. This formulation uses a transfomation that treats the two gauge fields in a quite different way from the beginning, while our procedure is more symmetric with respect to the two fields. We report the results and compare against our formulation. The two procedures yield different linear combinations for the fields after diagonalization in terms of the original fields, as expected, however results, such as the mass of the final massive gauge field turn out to agree completely. Thus the two methods appear to be consistent.
We should note that our model does not permit a distinction of the dark coupling to the hypercharge or the electromagnetic To obtain this possibility, we have to use a model based on This is deferred to the future. For the time being we examine the two models B and C. The former model B imitates the case where only the dark sector has a spontaneous symmetry breaking, and we study the effects of the symmetry breaking on the remaining unbroken sector. This is rather tricky, in the sense that the diagonalization has to be done numerically. The latter model C imitates the case where both sectors couple to the scalar field with equal strength. Spontaneous symmetry breaking is symmetric and the diagonalization is straightforward.
The formulation on a space-time lattice appears in section 3, where we take up the diagonalization procedure in this context. Section 4 contains the determination of the higgs phase transition in the lattice formulation for the three models. In particular it contains the Wilson loops and fine structure constants of all models in the symmetric phase. Then the broken phase for model B is described, where the massive and massless gauge fields yield Yukawa mass parameters and fine structure constants respectively. The phase diagram of model C comes next, followed by the plaquettes and correlators of various fields versus as well as their correlators.
Section 5 contains our conclusions.
2 The model in the continuum formulation
The full model in the continuum reads:
where is the Maxwell tensor of
2.1 Gauge fields
The gauge sector of the model may be written in the form:
This bilinear form can be made diagonal and yield the alternative expression:
where
One should rescale the fields, to get kinetic terms of the standard form:
so that the Lagrangian density takes the form:
The first step towards diagonalization of the symmetric phase theory involves an orthogonal transformation, however the need to obtain canonical kinetic terms imposes a non-orthogonal factor. Thus the full transformation is the product of the two transformations, that is:
or:
It is obvious that the transformation matrix is not orthogonal. We may also write down the inverse relation:
2.2 Matter fields
Let us next examine the coupling to the scalar fields, in particular the covariant derivative:
The dimensionless quantities represent the charge of the scalar fields in units of the respective couplings The part which will give the mass terms reads:
The mass of the scalar field derives from the scalar potential
which has been introduced in equation (1). Minimization of this expreszion yields the equality:
We may rewrite the above expression in terms of the primed fields and set so that the mass term will read:
After diagonalization we get:
where:
In the following, in the interest of simplicity, we set The mass term for the gauge fields reads then:
We may express the final potentials, and in terms of the original potentials, and via the formulae:
The field is massless, while field is massive. We also write down the inverse relations:
Model A:
For the values pertaining to model A, we get:
We may also express the and fields in terms of and
The interaction Lagrangian density for fermions may be written in the form:
Of course:
Model B:
For the values pertaining to model B, we get:
The inverse relations read:
The mass eigenvalue equals:
The interaction Lagrangian density for fermions may be written in the form:
Model C:
For the values pertaining to model C, we get:
We may also express the and fields in terms of and
The interaction Lagrangian density for fermions may be written in the form:
2.3 Alternative approach
The approach followed, e.g. in reference [27], is somewhat different from ours. In the following we examine the gauge sector. In [27] the authors start with the fields and rotate towards via a specific matrix. We will derive this matrix from the beginning, under the demand that should be just a multiple of with no admixture of
We substitute it into the expression:
This expression should reduce to
so one gets the relations:
One may solve this system (with the additional information that, for The result reads:
that is:
Considering the gauge part of the action, for the case:
To summarize, the diagonalization procedure is performed through the change of variables:
with their inverses:
After substitution one gets:
Following the symmetry breaking, setting for simplicity, the term
will contribute and we must diagonalize the total action, given by:
The covariant derivative may be rewritten in the form:
Thus the mass term resulting from this form reads:
One may diagonalize the matrix through the transformation:
Thus the new fields read:
so that:
By inspection of equation (53) and (54) it appears that the field is massless, while the is massive. The inverse relations of equations (55) read:
The non-zero eigenvalue for the gauge fields equals:
The fermionic lagrangian density reads:
We now continue with the special cases.
Model A:
and:
We remark that:
so the appropriate fields and to deal with after symmetry breaking equal and The fermionic lagrangian density reads:
Model B:
so that:
In this case the appropriate fields are approximately and The approximation is exact for
The fermionic lagrangian density reads:
Model C:
so that:
Once more:
so the appropriate fields and to deal with after symmetry breaking equal and
The fermionic lagrangian density reads:
3 Formulation on a space-time lattice
The model in the continuum reads:
where is the field strength tensor of the gauge fields. The non-compact Euclidean lattice action of the 4-dimensional dark matter model reads:
The quantities and are numbers which determine the interaction of the scalar field with the two gauge fields. In this paper we will consider the cases: model A, where model B, where and model C, where
3.1 Mixing
The gauge part of the Euclidean action for two kinetically coupled gauge fields may be transcribed from the continuum to the lattice form as follows:
where:
This expression may be diagonalized and take the form:
or:
The eigenvalues read:
The orthogonal transformation from the field strengths and to the new ones, and is a very complicated expression. In this work we will concentrate in the case
where the eigenvalues simplify to:
and the transformation is much simpler:
equivalently:
Let us remark at this point that equations (80) for the quantities and may also contain an overall minus sign. Of course, the potentials and will transform to and in a quite similar way:
equivalently:
We also note the definitions:
The new fields are massless just like the old ones.
We now consider interaction with a scalar field, the covariant derivative reading:
Notice that the initial expressions e.g. have been replaced by since we use the lattice version for the potentials. The quantities and are pure numbers which determine the strength of the interaction of the gauge potentials and with the scalar field. We write more compactly:
Thus we expect that, in the symmetric phase, we will have two massless fields, with the new charges reading and
We recall that in model C we set so that and As we remarked previously, see (63) and (70),
We expect that the appropriate fields to consider is which, in the broken phase, will develop a mass (since it is proportional to which is massive) and which will remain massless, since it is a multiple of the massless field
In model B we set so that and as is evident from equations (88). In this case, the appropriate basis to consider for is the one consisting of the massive and the massless However, for the appropriate basis is a combination of and depending on the value of
4 Results
4.1 Higgs phase transition
We examine the phase transition between the symmetric and the broken phase for We also set: In this work we will examine the three different models mentioned above, differing in the values of and We rewrite the lattice action substituting:
Notice that:
so that:
Model A is defined by In this case so that the scalar field interacts only with the field and it is also convenient to express the gauge kinetic terms as functions of and :
One may now perform a rescaling of the gauge fields: and end up with:
It is apparent that this action describes a non-interacting gauge field along with a field coupled to the scalar field. The lattice gauge couplings and are times the nominal ones. Thus we expect that the phase transition lines in the plane will differ significantly from the cases B and C, where we will see that no rescaling is necessary.
We examine model B, defined by for which:
This model describes a non-interacting gauge field along with a field coupled to the scalar field.
Model C is defined by In this case and no rescaling is needed:
It is obvious that model C shares some features with model B; the scalar field couples only with field (in model B) and with field in model C. Despite the difference in the choice of gauge fields, the mixing coupling stays smaller than as it will be observed in the following, for instance in figure 4 below. As towards various quantities get increasingly large values.
In the numerical study we examine the quantities
We also define a new variable considering the part of the action referring to the coupling of the scalar to the gauge fields:
where we have set:
The definition of the new, angular, variable is:
Initially we performed hysteresis loops for constant and varying to locate the phase transition points. One such loop is depicted in figure 1, which involves the loop for the gauge field in model B and, in figure 2, the loop for the angle of the scalar field in model B. In both cases It is remarkable that, in figure 1, the quantity does not change at all, since the does not couple to the scalar field and is not influenced by the spontaneous symmetry breaking. For the quantity the loop indicates a first order phase transition; the same conclusion is valid for the quantity depicted in figure 2. These statements are confirmed by long runs, which permit more accurate determination of the critical points and which clearly display two-state signals.
In figure 3 we depict the phase diagrams for using a lattice. It turns out that the curve for model A is very different from those for models B and C, while the latter two models almost coincide.
We depict, in figure 4, the phase diagram for all three models at using once more a lattice. Of course, the phase diagram makes no sense for We see that the phase transition lines for model C lies at higher values of than model B, while model A lies again systematically at larger In the latter case, model B is represented by the lower curve. It is apparent that it is possible for a given to select a value for which places model B in the broken phase and model C in the symmetric phase. In particular, for models B and C may lie in the broken phase, but, for there exist values for and for which model C lies in the symmetric phase, although model B may still lie in the broken phase. This conclusion follows by inspection of figures 3 and 4. In the following we concentrate on comparing B against C, since model A differs significantly from B and C, even for
4.2 Symmetric phase, dependence, models B and C
4.2.1 Decoupling
It is necessary to know the correlators of the various fields, so that we may calculate expectation values of fields decoupled from each other. Tree-level calculations have been described in section 3.1. The relevant correlators are given by the expressions:
The results are shown in figure 5. The correlators for all models are non-zero and they diverge in the limit. The important result is that the correlator takes invariably a value compatible with zero, which was the point of transforming from and to and in section 3.1. It appears that the transformation to and has been successful in decoupling the fields in the symmetric phase for all models.
Notice that
Since we predict that in the symmetric phase, a prediction that we have verified numerically. Notice the fact that, although we simulate the model using the and fields, the simulation automatically performs the necessary diagonalization and yields zero correlation between the and fields.
4.2.2 Preliminary calculation of a single plaquette
Another piece of information that will be useful is the expectation value of a model with just one gauge field. The action reads:
The results are depicted in figure 6 and it turns out that the curve may be approximately represented by the expression
for lattice volume The important point is that the plaquette varies as the inverse of a fact that will be important for us in the sequel. The parameter gets smaller values as we increase the lattice volume.
4.2.3 dependence of the decoupled fields in the symmetric phase
Now we are in a position to proceed with the two-field models with a kinetic coupling term. The picture of the symmetric phase is similar for the two models, since the symmetry breaking, which discriminates between them is absent. The parameters read: and We just consider The quantities and are shown in figure 7, for the symmetric phase for models B and C (where ). The two quantities are represented by the middle curve and they are equal (this agrees with the prediction based on the vanishing of the correlators, already mentioned previously, equation (107)). The quantities and start from moderate values and increase with However, we recall that the and fields are highly correlated, due to the kinetic mixing term, so one should rather consider the fields and which are uncorrelated in the symmetric phase, as seen above. Since the two fields are uncorrelated, we expect to get results which will behave as suggested by (109), their only difference being that is inversely proportional to while is inversely proportional to In figure 7 one may see the results for (it is the uppermost curve) and we may inspect the divergence of this quantity as On the other hand, (the lowest curve) decreases with increasing since never vanishes.
We now proceed with fitting the expectation values of and with and respectively. The fits yield:
We observe that the numerators are somewhat different from found from the single gauge field before, equation (109). It means that the system, although it is similar to a set of independent gauge fields it is not entirely the same, since interaction with the scalar field is also involved.
We have also calculated the Wilson loops for and and we have fit them using the expressions:
The results are shown in table 1.
We may calculate the tree-level fine structure constant using the relations and If we eliminate among the two expressions, we get One may compare the value of from the Wilson loops against the tree-level prediction, after substituting with the quantities since these are the expected couplings for
We observe that the fine structure constants calculated from the Wilson loops are in general larger than the tree level values. However, the fine structure constants follow the variation of the tree level parameters. The coupling increases, while decreases with
We conclude that differs from just because the two fields and depend on and respectively. There is no important difference between models B and C in the symmetric phase. One may fit the data and discover that:
Table 1: Predictions and results for all models in the symmetric phase.
4.3 Broken Phase: Model B
We set which places the system in the broken phase. We start with the dependence of various quantities on . Figure 8 shows that neither nor are appropriate for the study of the model, since both correlators and are different from zero (with the exception of at There is a striking feature of these diagrams that needs explanation, namely the fact that the correlators appear to take very large values for We shall deal with this later on. We have restricted plotting to since the values of for larger are too large to depict in the graph. Another remarkable feature in the graph is the fact that the correlator vanishes for We will also come to this later on.
We calculated the plaquettes, which will prove useful in the sequel. The results for and (figure 9), are not equal any more: the plaquette of the visible field starts from a fairly large value, since it is in the unbroken phase, while the dark field starts at small value, since it lies in the broken phase; both increase with These results are consistent with the ones in figure 1, where the and the loops are shown. On the other hand, increases, while slightly decreases with It appears that the pair behaves as expected from the tree level calculations: takes a relatively large value, since it corresponds to an unbroken component, while takes small values, compatible with a broken phase observable. On the other hand the plaquette values, with the interesting exception of are also taking very large values for much like the correlators in figure 8.
To understand the abrupt change in the behaviours of the correlators in figure 8, as well as of the and plaquettes in figure 9, we depict in figure 10 the phase diagram for the two values and If we concentrate in i.e. the values used throughout this work, it is seen that the system is in the broken phase for (lower curve), and in the symmetric phase for (upper curve). Thus the large values seen before in the plaquettes are connected with the passage of the system to the symmetric phase. The same can be said about the correlator, which vanishes abruptly for this may also be attributed to the passage to the symmetric phase. In addition, the plaquette takes the symmetric phase value in this parameter region.
As explained above, the problem with the results so far is that the field correlators as well as don’t vanish, that is the corresponding fields exhibit a non-trivial mixing. It is desirable to find linear combinations of the fields, which yield uncorrelated fields. We rotate the fields to through an angle
and vary this angle until the correlators of the resulting fields vanish. The fields and are the lattice analogs of the fields and mentioned in the continuum treatment of the model. Some interesting special values for are:
(a) for which At the and fields are exactly equal to and respectively, while for these expressions also involve admixtures of and respectively (see table 2).
(b) for which that is the linear combinations appropriate for models A and C.
For small we expect that will be approximately equal to as we saw in our analysis in the paragraph about matter fields. By the same analysis, will be proportional to (equation (28) of subsection 2.2). The relevant angles for selected values of are shown in table 2. One may see that the appropriate fields, and are not exactly equal to and as the tree level considerations suggest, however they are mainly (with small admixture) and (with small admixture) for a quite large range of say up to For bigger the contributions of in and the one of in become significant.
In figure 11 we show the results of the simulation for and versus It appears that is approximately constant (slightly decreasing), while increases slightly as increases. This is to be expected, since equals mostly which does not couple to the scalar field, while equals mostly so that it is influenced very much by the symmetry breaking. As increases and tends to approaches while approaches the fields that mostly characterize the symmetric phase. This transition is shown in figure 11 for around 1.75.
Table 2: Rotation angles for the construction of uncorrelated fields in the broken phase of model B.
Table 3: Fit parameters for the fields and as well as the uncorrelated fields and in the broken phase of model B.
We have also calculated the Wilson loops for and and fit them using the expressions:
The results are shown in table 3. In fact, we have proceeded up to (not shown in the table) and found that is compatible with zero, which agrees with the previous conclusion that the system passes to the symmetric phase for these values of The value of does not change much. It is interesting that is approximately constant and equal to which is expected, since this field is presumably not coupled to the scalar field.
We see that, for coincides with the field, which interacts with the scalar field and is best simulated with a Yukawa form for the potential in the broken phase; on the contrary, for again, coincides with the field, which does not interact with the scalar field, so one should expect the behaviour appropriate to pure gauge fields. For non-zero there is a difference, in the sense that it is necessary to include a slight admixture of into apart from the This is not predicted by the tree level considerations. Let us recall equations (28):
The tree-level field corresponds to while corresponds to We find that, for small derives from only, and this is consistent with the results of table 2, where the corresponding field is basically (with an admixture of however). On the lattice we need an admixture of and to get the uncorrelated fields and Thus, for small it turns out that is basically and yields a Coulomb potential, while is basically and yields a Yukawa potential. It is reasonable to expect Coulomb expressions for the fields and Yukawa expressions for the fields; the numerical results confirm these expectations. Having fixed the linear transformations for each value of we calculate the potentials for and and fit with a Yukawa form with mass and with a Coulomb form, where the fine structure is
We have also reported, in table 3, the masses of the scalar field, to compare with the mass of the gauge field. It turns out that the mass of the scalar field is approximately independent from while is comparable to for small but it drops substantially when increases. The scalar field can possibly decay into the gauge and other fields, while the products of the decay may not include the scalar, especially for
4.4 Broken Phase: Model C
We start our study of model C with its phase diagram in figure 12. In this figure we show the phase diagrams for model C at six values of namely and If we focus on the value used throughout our paper, we observe that, for the system lies in the broken phase while, for it lies in the symmetric phase. The phase transition point, when is estimated to be around Thus the system is estimated to move from the broken to the symmetric phase when increases, while the remaining parameters are kept fixed. This means that big push the system towards the symmetric phase.
In fact, for phenomenological studies of dark matter, should be restricted to small values. We examine what will happen for other values of the gauge coupling Consider, for example the two middle curves in figure 12, corresponding to and If we concentrate in we see that places the system in the broken phase, while places it in the symmetric phase. The estimated critical value for is around 0.67. Finally we consider the two curves on the left, corresponding to and We may now easily see that the values take us deep in the broken phase for and in the symmetric phase for with an estimated critical value around There follows the conclusion that critical values may be small, provided is small enough.
A distinctive feature of the phase diagram worth mentioning is the behaviour of the phase transition lines near The coupling may not be smaller than For instance, for (the rightmost curve), the phase transition line moves to smaller and larger however, for large the line becomes vertical at the value since it may not become smaller that In section 3.1 we have seen that, when and as is the case in model C, it is appropriate to describe the system through the massive field (which couples with the scalar field via the dimensionless charges ) and the massless field, which does not couple to the scalar field, since According to equation (79) the couplings for these fields in the symmetric phase equal: In the symmetric phase, it is evident that should be positive, meaning that may not be smaller than It appears that this feature insists in the phase diagram, although we examine the broken phase and we have not done any transformation to the and fields.
The line has an abrupt change and finally it assumes its final vertical form for or larger. A rather unexpected result is that the value for which the vertical part starts, is the same (within errors) for all values.
We now go ahead with the study of the dependence of the quantities and on (), which may be seen in figure 13. The variation of the various quantities is fairly mild, up to where there is an abrupt change. This behaviour can be explained with the help of the phase diagram of figure 12. The quantities and in figure 13 are equal, since their couplings to the scalar field are equal. However, we have found that the correlators cannot be neglected, so one should diagonalize the action, using the and fields, for which the correlator vanishes. The field is the one that couples to the scalar field, so that the quantity takes on rather small values. On the other hand, does not couple to the scalar field, so the values of are comparatively large; they actually decrease slightly with increasing since they are expected to be proportional to These explanations hold up to At that point, as explained previously, the system passes over to the symmetric phase, so the field is now proportional to and takes very large values. This behaviour is also seen in and since they may be expressed as functions of On the contrary, for the field nothing changes and it decreases for as well.
In table 4 the mass of and the fine structure constant of as well as the mass of the scalar particle for model C may be found. The fine structure constant decreases with as expected. The mass of also decreases. Actually, if one continues with the simulations for even larger this mass will become compatible with zero and the system will pass over to the symmetric phase after On the other hand, for small the mass of is larger than the scalar mass[30]; but this is reversed for and it is expected that the gauge field cannot give scalars through its decay, if grows large enough. In addition, the coupling in table 4 is approximately equal to in table 1, that is, the corresponding field in the symmetric phase. Thus the field in the symmetric phase, which is not coupled with the scalar field does not change essentially after symmetry breaking.
Table 4: The parameters and in the broken phase of model C.
5 Conclusions
We have studied a model, suitable for dark matter inquiries, among others. The model is based on a product group, with factors, and we have identified the resulting gauge groups in a couple of models with varying scalar field charges. The model is more general than its motivation.
(1) The tree-level mass predictions in the continuum [25] for the massive gauge fields predicted by the usual method equal the predictions of the method proposed in this work. One may see this readily by comparing (24) against (60) for model A, (30) against (65) for model B, and (33) against (68) for model C. That these predictions are identical lends confidence that the approach proposed here is on the right track.
(2) The tree-level mass predictions in the continuum behave quite differently from the corresponding lattice results: they tend to diverge, when approaches equivalently when approaches one. On the contrary, their lattice counterparts approach zero in this limit, as may be seen inspecting the column in table 3 and the the column in table 4. The source of this strong disagreement may be traced to the fact that lattice calculations are non-perturbative, so they can, e.g., detect phase transitions, while tree-level calculations cannot. In our case, increasing to approach moves the system towards a phase transition, which cannot be seen from the tree-level calculations.
(3) The fine structure constants displayed e.g. in tables 1, 3 and 4, are more or less the same for small enough , despite the fact that they refer to different situations (for example, symmetric versus broken phases). In the symmetric phase the quantity increases, while decreases with increasing in qualitative agreement with the tree level predictions In the broken phase of model B the quantity which corresponds to (this does not couple to the scalar field), is essentially constant. This is in sharp contrast with the behaviour of the fine structure constants in the symmetric phase. On the other hand, in the broken phase of model C the quantity decreases with increasing This is qualitatively similar with its behaviour in the symmetric phase. On the contrary the field yields a Yukawa behaviour.
(4) The masses in model B of the gauge field which interacts with the scalar field, do not depend appreciably on while the masses of do so. These masses are equal for and then the mass of the gauge field becomes smaller, as increases. The masses of the scalar field in model B slightly increases with and becomes bigger than the mass of the gauge field In model C the mass of the gauge field is bigger than the scalar mass for but it diminishes as increases and becomes smaller than the mass of the scalar field at The tree-level prediction for the scalar field mass is a value that does not depend on for either model.
(5) The system may move from the broken to the symmetric phase, when is changed. According to figure 10, there is a phase transition from the broken to the symmetric phase in model B, as becomes bigger than 1.75. This explains the qualitatively different behaviour for in figures 8 and 9. Similarly, according to figure 12, there is a phase transition from the broken to the symmetric phase in model C, as becomes bigger than 1.63. This explains the qualitatively different behaviour for in figure 13. The critical values of move to smaller values (for instance 0.67 or 0.18, as depicted in figure 12) if is chosen to be smaller, i.e. for stronger gauge coupling.
(6) We have performed simulations of the model for negative and the results for the mean values and are equal to the mean values for
Future work may include simulation of a model based on to provide the possibility to study the main physically different and realistic couplings of the dark to either the hypercharge or the electromagnetic
Acknowledgements
We would like to thank the administration of the ARIS Computer Center for computer time allocated to us, as well as Professor A. Tsapalis for help with the code parallelization. Thanks are also due to Dr P.Dimopoulos for illuminating discussions, as well as to professor N.E.Mavromatos for enlightening discussions and for reading the manuscript.

### v2

@@
 for stronger gauge coupling.
+In an attempt to relate lattice parameters with continuum quantities we also write down the effective potential in the continuum at one loop, where the gauge coupling should be replaced by with a quantity independent from This form describes the effective coupling in the presence of kinetic coupling, as explained in the paper; it is expected to provide valuable non-perturbative information.
+The result is that the characteristic of two minima persists, as one increases however the separation of the two minima becomes very small.
+This means that the phase transition towards the symmetric phase is missed by the one-loop effective potential, but it appears that the system will possibly move to the symmetric phase near if higher loop corrections are included; we are working on this.
 (6) We have performed simulations of the model for negative and the results for the mean values and are equal to the mean values for
-Future work may include simulation of a model based on to provide the possibility to study the main physically different and realistic couplings of the dark to either the hypercharge or the electromagnetic
+Future work may include:
+(a) Numerical computation for the functions for the various couplings via a non-perturbative calculation.
+111We thank the referee for raising this issue.
+To this end one might compute the dependence of our results on the lattice size, to get an estimate of the dependence on the lattice spacing and the approach to the thermodynamic limit.
+This calculation would allow for a connection with physical quantities and we postpone it for a separate publication.
+(b) Simulation of a model based on to provide the possibility to study the main physically different and realistic couplings of the dark to either the hypercharge or the electromagnetic
 Acknowledgements